1) Find the critical value za/2 that corresponds to a 95% confidence level. Explain how to enter
this in your calculator
Answer #1

Solution:

We are given

c = 95% = 0.95

α = 1 - c = 1 - 0.95 = 0.05

α/2 = 0.05/2 = 0.025

Z α/2 = 1.96

(by using Ti-84/83 calculator)

[Use 2ND > DISTR> Select 3:invNorm( function for finding above value]